OpenStudy (anonymous):

converting 54 degrees farenheit to celcius, i subtracted 32 what do i do know ?

OpenStudy (anonymous):

F = C * 9/5 + 32 C = (F - 32) * 5/9 You can check (and derive) these formulas if you remember that water freezes at 0 C and 32 F, and boils at 100 C and 212 F.

OpenStudy (anonymous):

btw its converting weather

OpenStudy (anonymous):

It doesn't matter what you're taking the temperature of. Temperature is temperature.